OpenStudy (anonymous):

If sinA=3/5, find sin4A (Dealing with half and double angle formulas)

OpenStudy (shamim):

sin4A=2*sin2A*cos2A=2*2*sinA*cosA*(cos^2A-sin^2A)

OpenStudy (anonymous):

if sinA=3/5 then first find cosA ? and proceed futher..(alterantely u can find cos2A using cos2A=1-2sin^2A and then cos4A using cos4A= 2cos^2(2A) -1 and then fin sin4A)

OpenStudy (shamim):

r u able to find out the value of cosA when sinA=3/5